find the two consecutive positive even numbers the sum of whose squares 340.se[arate the two numbers with a comma
Answer by checkley77(12844)

X^2+(X+2)^2=340
X^2+X^2+4X+4=340
2X^2+4X+4-340=0
2X^2+4X-336=0
2(X^2+2X-168)=0
2(X+14)(X-12)=0
X-12=0
X=12 ANS.
12+2=14 ANS.
(12,14)
PROOF:
12^2+14^2=340
144+196=340
340=340